OpenStudy (anonymous):

Input in standard form the equation of the given line. The line with m = -3/2 and b = 6

OpenStudy (texaschic101):

In y = mx + b form, the m position is the slope and the b position is the y intercept. So we have : y = -3/2x + 6 But we want standard form (Ax + By = C) y = -3/2x + 6 -- add 3/2x to both sides 3/2x + y = 6 -- can't have fractions, so multiply equation by 2 3x + 2y = 12 ===> standard form

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

no. Thanks. Can you help me with one more?

OpenStudy (texaschic101):

I can try

OpenStudy (anonymous):

Input in standard form the equation of the given line. The line with m = -1 /2 and b = 1

OpenStudy (texaschic101):

y = -1/2x + 1 This is not the answer, but do you know how I arrived at this ?

OpenStudy (texaschic101):

when m is -1/2 and b is 1, your equation becomes this. Do you understand so far ?

OpenStudy (anonymous):

yes

OpenStudy (texaschic101):

y = -1/2x + 1 --- add 1/2x to both sides 1/2x + y = 1 -- no fractions, so multiply entire equation by 2 x + 2y = 2 ==> standard form

OpenStudy (anonymous):

thanks

OpenStudy (texaschic101):

sure thing :)